OpenStudy (anonymous):

A line passes through the point (−8,-3) and has a slope of 3/2. Write an equation in slope-intercept form for this line.

OpenStudy (panda.exe):

hi

OpenStudy (anonymous):

@hockeychick23 that's wrong

OpenStudy (anonymous):

Well the slope intercept form is y = mx + b

OpenStudy (anonymous):

m being slope, x being the x value, and b being the y intercept.

OpenStudy (anonymous):

So plug in what we know y = -3 so -3 = mx + b x = -8 so -3 = m(-8) + b The slope, m, is 3/2 so -3 = 3/2(-8) + b Now simplify

OpenStudy (anonymous):

-3 = -12 + b +12 +12 9 = b

OpenStudy (anonymous):

so the final equation is y = 3/2x + 9

OpenStudy (anonymous):

@ryryfofo

OpenStudy (anonymous):

Any other questions?

OpenStudy (anonymous):

@jackmullen55 nope.. thanks!